OpenStudy (anonymous):

Find the value of y when x equals -4 in this equation. 5x + (-3y) = 13 A. 9 B. -11 C. -5 D. 5

OpenStudy (anonymous):

b -11

OpenStudy (anonymous):

plz just dont give away the answers 1

OpenStudy (yrelhan4):

You just need put the value of -4 in place of x and solve for y.

OpenStudy (anonymous):

5(-4) + (-3y) = 13 -20 + -3y = 13 -3y = 13 + 20 -3y = 33 3y = -11

OpenStudy (anonymous):

*y = -11

OpenStudy (anonymous):

put x=-4 in the equation 5(-4)-3y=13 -20-3y=13 -3y=13+20 -3y=33 y=33/-3 y=-11